The total value of the pennies and nickles was $14.50.
Emet and callaway counted the coins and found that the total value was 450 coins. How many coins of each type did they have?

Answer by Alan3354(69443)

Here's a similar problem solved.
*****************************************************
I have 20 coins. All of them are either one pound or 50 pence coins. They are worth £11.00. How many pound coins and 50 pence coins do I have. (Algebra)
--------
Thx for telling us it's Algebra, we might not have known that.
----
p = pound coins
f = 1/2 pound coins
---
p + f = 20
p + f/2 = 11
---
I would solve by elimination (an Algebra thing).